OpenStudy (anonymous):

If F(x) = 3x + 2 and G(x) = x2 + 4, what is F(G(x))?

OpenStudy (rea201):

F(G(x))=3(x^2+4)+2 F(G(x))=3x^2+12+2 F(G(x))=3x^2+14